# Pay Attention to Your Neighbours: Training-Free Open-Vocabulary Semantic Segmentation

### v1

@@ Abstract
Keywords:
Semantic segmentation Open-vocabulary Training-free
@@ 1 Introduction
Open-vocabulary semantic segmentation (OVSS) has emerged as an appealing alternative to traditional close-set approaches, as it can handle novel categories that may not have been seen during training. Fueled by the zero-shot performance of vision-language models in visual recognition tasks, most recent OVSS approaches [2, 14, 19, 37, 39, 40, 46, 28, 38, 22, 39] are inspired by the Contrastive Language-Image Pre-training (CLIP) [27] paradigm. A popular family of approaches integrates a fully-supervised training step [2, 14, 19, 37, 39, 40, 26], where pixel-wise masks from a limited set of categories are leveraged to transfer language-visual alignment at image level to pixel granularity. The pixel-level annotation assumption can be relaxed by considering a weakly-supervised adaptation dataset, where only image-text pairs are accessible [8, 6, 43, 34, 7]. These methods, however, still require a substantially large set of either pixel-wise or image-text annotations, which typically present a large overlap with the open-set categories in the test datasets. Furthermore, the final model performance may be biased towards the training dataset selected for adaptation. It is noteworthy to highlight that COCO-Stuff is typically employed in fully-supervised adaptation, which contains 171 unique annotated classes with significantly high overlap with datasets used for evaluation (please see Table 1 in [39]).
@@
To meet the requirements of real-world applications, where access to large labeled datasets is rare and one cannot anticipate novel classes, we focus in this work on training-free, a more challenging and realistic scenario without access to additional data for adaptation. Due to their practical relevance, there exists a growing literature on these approaches [46, 28, 35, 16, 32, 4, 3]. Nevertheless, we found that despite their categorization as training-free strategies, most methods present important practical limitations, including: use of additional pre-trained networks, such as MoCov2 or DeiT, or unsupervised object localization approaches trained on additional datasets [28, 35, 23]; requiring pre-trained text-conditional generative models, e.g., stable diffusion, to generate multiple additional images for novel concepts [3, 16]; tuning many hyperparameters [3, 23], whose values change across datasets [3]; or integrating a weakly supervised reward function for hyperparameter tuning [23], among others. Thus, devising novel alternatives that can relax these requirements and present simple solutions in a more realistic training-free manner is paramount to deploying these methods in real-world problems.
@@
In dense prediction tasks such as semantic segmentation, the crucial aspect of localization is often overlooked when employing vision transformers (ViTs) [10]. Notably, ViTs emphasize global representations, particularly through their [CLS] token, which leads to suboptimal dense prediction performance. Therefore, we advocate that devising novel and effective methodologies hinges on considering design choices that favour segmentation. The recent concurrent work in [32] shares a similar line of research, and investigates the inherent problems of CLIP in dense prediction, proposing minor adjustments over the baseline model. In particular, the authors identified that the baseline’s poor performance is caused by a spatial misalignment of the patch representations, pointing to CLIP’s self-attention modules as the source of the problem. More concretely, they argue that CLIP learns spatial-invariant visual features, beneficial only in image-level tasks. To address this limitation, a new self-attention mechanism was introduced in [32], which re-organizes the spatial information. Nevertheless, despite its remarkable performance compared to CLIP, it does not guarantee semantic correlations across local tokens, as it does not have an explicit mechanism to attend to each token’s neighbours and therefore ensure spatial consistency. As depicted in Fig. 1, although SCLIP [32] captures semantic context better than the standard attention in CLIP, it yields unstable attention maps on nearby tokens.
@@
Motivated by these limitations, in this work we do due diligence in the task of training-free OVSS and propose a straightforward baseline with minimal modifications to vanilla CLIP. The implemented changes are motivated by the specificity of dense prediction problems, which have been mostly disregarded in existing methods. Specifically, we identify the potential limitations of CLIP in image segmentation and modify its visual encoder, particularly in the final layer, to enhance its localization capabilities. Furthermore, we analyze the concurrent work in [32], exposing that despite enhancing the context localization of the vanilla CLIP, it suffers from unstable attention maps. We argue that this is due to the lack of a proper mechanism that ensures spatial consistency, and propose a simple solution that encourages each patch to attend to its neighbours, generating consistent attention maps across adjacent patches. As demonstrated empirically through comprehensive experiments on 8 popular OVSS benchmarks, our strong baseline Neighbour-Aware CLIP (NACLIP), achieves state-of-the-art performance without requiring additional data, auxiliary pre-trained networks, or extensive fine-tuning of hyperparameters.
Figure 1: Attention maps of the final layer. For each of the patches shaded in red (denoted with (a) to (d)), the attention maps of the final visual encoder layer are presented for vanilla CLIP [27], SCLIP [32], and our method. We have identified two problematic phenomena in the attention maps of CLIP and SCLIP, both stemming from a lack of mechanisms to properly attend to patches’ neighbourhoods. First, as depicted in (a) and (b), attention intensity is sometimes dispersed among distant patches, neglecting the vicinity of a patch. Additionally, adjacent or closely located patches sharing the same real-world category and even similar visual characteristics can have inconsistent attention maps. For instance, while SCLIP generates a quality attention map for patch (d), its attention map for (c) is notably different and fails to focus on the desired object. By explicitly promoting attention to neighbours, our method produces consistent attention maps across adjacent patches.
@@ 2.1 Adapting Large-Scale Vision-Language Models
The field of machine learning, and more specifically deep learning, is undergoing a learning paradigm shift with the emergence of vision-language models trained at a large scale. In particular, the Contrastive Language-Image Pre-Training (CLIP) method [27] has shown unprecedented success, mainly due to its remarkable zero-shot and few-shot transfer capabilities on visual recognition tasks, notably in the context of classification. Fueled by its generalization and transfer capabilities, a realm of approaches have arisen to either improve its zero-shot performance by modifying its training on image-text pairs [41, 33, 12], or efficiently adapt to novel tasks with only a few labeled images [44, 13, 42, 30, 15, 29]. Nonetheless, the pre-training is performed at the image level, and therefore typically only the class token is trained to capture the global information, which limits the applicability of these approaches for dense prediction tasks.
@@ 2.2 Open-Vocabulary Semantic Segmentation
The outstanding transferability of CLIP in classification has surged to a rapidly growing literature in open-vocabulary semantic segmentation (OVSS) [2, 14, 19, 37, 39, 40, 46, 28, 38, 22, 39], which attempts to segment novel concepts in a given image without explicit supervision for them. These methods can be categorized into three main groups, depending on the level of auxiliary data required for adaptation: fully-supervised, weakly-supervised, and training-free. Fully-supervised OVSS approaches [2, 14, 19, 37, 39, 40, 26] adapt the pre-trained CLIP to semantic segmentation by exploiting a large labeled set containing pixel-wise annotations from a set of classes. The main idea of these approaches is that, by leveraging a large pixel-labeled dataset of an arbitrary limited set of categories, the model can learn to perform semantic segmentation, while maintaining the remarkable transfer capability of CLIP. Nevertheless, the adaptation datasets employed in this family of approaches typically present a high class overlap with the open-set of testing images. Instead of requiring pixel-wise masks, weakly-supervised OVSS resorts to additional datasets with image-level tags, typically in the form of image-text pairs [8, 6, 43, 34, 36, 21, 7]. A common strategy is to use these large image-text pair sets as supervision during adaptation, where the categories present in each image are included in the text, e.g., “A racer rides a bicycle in the air”. The alignment between the visual and textual information is typically performed via a contrastive loss [36], similar to the pre-training of CLIP, which can be further enhanced by integrating additional learning strategies, such as online pixel clustering [21] or multiple contrastive losses between the features from original images and those restored from corrupted versions [6], for example. Nevertheless, a substantially large dataset is still required for adaptation, and classes appearing in each image need to be known in advance, presenting impractical considerations for real-world problems.
@@
Given the points raised above, we focus in this work on training-free OVSS [46, 28, 35, 16, 32, 4, 3, 18] where, ideally, access to additional data for adaptation is not allowed. Nevertheless, as stated earlier, most of these methods rather rely on leveraging auxiliary models pre-trained on large-scale datasets (e.g., ViT [28, 35, 23] or stable diffusion [3, 16]), or containing multiple components [3, 23], whose hyperparameters must be tuned, sometimes in a per-dataset manner [3], or using a validation set [23]. A different line of work attempts to improve the potential of ViTs in extracting dense visual features by modifying ViT’s self-attention [46, 4, 18]. This can be done, for instance, by directly using the value vectors [46], or by introducing additional pathways computing self-self attentions in parallel to the encoder blocks to aggregate output of several layers using the residual connections [18, 4]. A main difference with these works, however, is that we introduce explicit local spatial consistency in the self-attention of the ViT, a concept that has been disregarded by the existing training-free OVSS literature.
@@
The concurrent approach SCLIP [32] proposes a self-attention mechanism encouraging each token to attend to itself and the positions sharing similar information. In particular, the authors state that an attention map with maximum values on the diagonal leads to proper localization. Nevertheless, since query and key vectors are not normalized in SCLIP, there is no guarantee that the maximum values of the attention map fall on the diagonal. Hence, if an outlier patch has a key vector with a magnitude much larger than the others, patches attend to the outlier more than themselves. Furthermore, even though in a well-localized model each patch should attend to itself with a high intensity, localization is not limited to self, and the vicinity of each patch should be taken into account as well. This is particularly important in segmentation, where adjacent patches often represent the same class. As apparent in our qualitative analysis (Fig. 1), despite SCLIP improving the spatial localization of attention maps over CLIP, it often generates contrasting attention maps for adjacent patches containing the same real-world object. This demonstrates that in this method, spatial localization in the neighbourhood of a patch does not typically go far beyond the patch itself. Please note that although for a given patch, SCLIP attends to the patches that share similar query or key vectors, neighbouring patches do not necessarily do. Besides, we argue that only attending to patches that share similar key or query vectors (and also self) is suboptimal, which is demonstrated empirically in Fig. 1. The proposed method can handle these limitations by adding an explicit mechanism that enforces patches to attend to their neighbours, imposing the local spatial consistency required for semantic segmentation.
@@ 3 Preliminaries
Problem Definition. In this work, we address the task of open-vocabulary semantic segmentation in a training-free scenario, i.e., with no supervision and without fine-tuning the parameters. Hence, for a given image $\mathbf{X}_{i}\in{\mathbb{R}}^{\Omega_{i}}$, with $\Omega_{i}$ denoting its spatial domain ($H\times W\times C$), and a set of arbitrary concepts described by natural language $\mathbf{t}_{j}\in\mathcal{T}$, our goal is to provide a segmentation mask per concepts present in the image.
@@ 3.1 Background on CLIP
CLIP [27] employs a joint training approach, aligning visual and text modalities into the same feature space. Since most CLIP-based OVSS methods utilize the vision transformer [10] version of CLIP, we focus on this architecture in what follows. In this model, the visual encoder comprises $L$ sequential blocks, each processing $1+hw$ tokens: the first one representing the [CLS] token, which captures global information, and the subsequent tokens each representing a patch. Thus, given an input image, it is initially partitioned into $hw$ non-overlapping patches ($h=H/P$ and $w=W/P$) each in ${\mathbb{R}}^{P\times P\times C}$, where $(P,P)$ denotes the resolution of each patch. Then, a linear transformation projects the sequence of $1+hw$ tokens from $C$ channels to a $D$-dimensional space, and positional embeddings are added to create the input for the first encoder block. The components of this model are described below. For the sake of simplicity in our formulation, we avoid flattening the 2D grid of patches.
3.1.1 Encoder Block.
Each encoder block $l$ receives a sequence of $1+hw$ tokens from the preceding block, $\mathbf{Z}^{(l-1)}=[\mathbf{z}_{\texttt{[CLS]}},\mathbf{z}_{11},\dots,\mathbf{z}_{hw}]$, and performs the following operations:
@@
In the equations above, $\operatorname{LN}$, $\operatorname{SA}$ and $\operatorname{MLP}$ denote layer normalization, the self-attention module, and the feed-forward neural network, respectively. Please note that skip connections refer to the addition of $\mathbf{Z}^{(l-1)}$ and $\mathbf{Z}^{\prime}$ in Eqs. 2 and 4. Furthermore, the combination of Eqs. 1 and 2 is commonly referred to as the self-attention block, whereas the combination of Eqs. 3 and 4 is called the feed-forward block.
3.1.2 Self-Attention Module.
Within the self-attention module, the sequence of token representations undergoes a linear transformation $\mathbf{W}^{qkv}$, yielding three sequences of $d$-dimensional vectors: query ($\mathbf{q}$), key ($\mathbf{k}$), and value ($\mathbf{v}$). Subsequently, a similarity measure is calculated between all the tokens. More specifically, for a given patch at position $(i,j)$ (and also for the first token corresponding to [CLS]) the dot product of $\mathbf{q}_{ij}$ and $\mathbf{k}_{mn}$ is computed for all $m\in\{1,2,\dots,h\}$ and $n\in\{1,2,\dots,w\}$. This measure is then scaled by $1/\sqrt{d}$ and passed through a $\mathrm{softmax}$ operation to calculate a weighted sum of $\mathbf{v}$ for the patch. Finally, the output is obtained via a projection using a linear transformation $\mathbf{W}^{o}$. Thus, formally, operation $\operatorname{SA}$ for the patch at position $(i,j)$ can be described as:
@@
Hereafter, we will refer to $\mathrm{softmax}\left(\text{sim}_{ij}\right)$ as the attention map of point $(i,j)$, and $\text{sim}_{ij}$ as its corresponding logits. Note that this operation is done in parallel for all the tokens, and while in practice a multi-head version of the self-attention is used, for simplicity, we have presented equations considering a single head.

### v2

@@ 1 Introduction
Open-vocabulary semantic segmentation (OVSS) has emerged as an appealing alternative to traditional close-set approaches, as it can handle novel categories that may not have been seen during training. Fueled by the zero-shot performance of vision-language models in visual recognition tasks, most recent OVSS approaches are inspired by the Contrastive Language-Image Pre-training (CLIP) [32] paradigm. A popular family of approaches integrates a fully-supervised training step [2, 16, 22, 44, 45, 46, 31], where pixel-wise masks from a limited set of categories are leveraged to transfer language-visual alignment at image level to pixel granularity. The pixel-level annotation assumption can be relaxed by considering a weakly-supervised adaptation dataset, where only image-text pairs are accessible [8, 6, 49, 40, 7, 42]. These methods, however, still require a substantially large set of annotations, which typically present a large overlap with the open-set categories in the test datasets (please see Table 1 in [45]). Furthermore, the final model’s performance may be biased towards the training dataset selected for adaptation.
@@
To meet the requirements of real-world applications, where access to large labeled datasets is rare and one cannot anticipate novel classes, we focus in this work on training-free, a more challenging and realistic scenario without access to additional data for adaptation. Due to their practical relevance, there exists a growing literature on these approaches [52, 33, 41, 19, 38, 4, 3, 21, 27, 11], which leverages CLIP [32] as the source of knowledge. However, some also utilize additional pre-trained networks like MoCov2 [17], DeiT [30], or unsupervised object localization approaches [35] trained on extra datasets [33, 41, 19], or pre-trained text-conditional generative models (e.g., stable diffusion) to generate multiple additional images for novel concepts [3, 19, 11]. We argue that methods using these auxiliary pre-trained models, having been trained on significant extra data and introducing additional weights, cannot be fairly compared to training-free methods that use only CLIP. Furthermore, some prior works present practical limitations such as tuning many hyperparameters [3, 27], whose values change across datasets [3], or integrating a weakly supervised reward function for hyperparameter tuning [27]. Thus, devising novel alternatives that relax these requirements and present simple solutions in a more realistic training-free manner is paramount to deploying these methods in real-world problems.
@@
In dense prediction tasks such as semantic segmentation, the crucial aspect of localization is often overlooked when employing vision transformers (ViTs) [12]. Notably, ViTs emphasize global representations, particularly through their [CLS] token, which leads to suboptimal dense prediction performance. Therefore, we advocate that devising novel and effective methodologies hinges on considering design choices that favour segmentation. The recent concurrent work, SCLIP [38], shares a similar line of research, and investigates the inherent problems of CLIP in dense prediction, proposing minor adjustments over the baseline model. In particular, the authors identified that the baseline’s poor performance is caused by a spatial misalignment of the patch representations, pointing to CLIP’s self-attention modules as the source of the problem. More concretely, they argue that CLIP learns spatial-invariant visual features, beneficial only in image-level tasks. To address this limitation, a new self-attention mechanism was introduced in [38], which re-organizes the spatial information. Nevertheless, despite its remarkable performance compared to CLIP, it does not guarantee semantic correlations across local tokens, as it does not have an explicit mechanism to attend to each token’s neighbours and therefore ensure spatial consistency. As depicted in Fig. 1, although SCLIP [38] captures semantic context better than the standard attention in CLIP, it yields unstable attention maps on nearby tokens.
@@
Motivated by these limitations, in this work, we do due diligence in training-free OVSS and propose a straightforward baseline with minimal modifications to vanilla CLIP. The implemented changes are motivated by the specificity of dense prediction problems, which have been mostly disregarded in existing methods. Specifically, we identify the potential limitations of CLIP in image segmentation and modify its visual encoder, particularly in the final layer, to enhance its localization capabilities. To ensure spatial consistency, we propose a simple solution that encourages each patch to attend to its neighbours, generating consistent attention maps across adjacent patches. As demonstrated empirically through comprehensive experiments on 8 popular OVSS benchmarks, our strong baseline Neighbour-Aware CLIP (NACLIP), achieves state-of-the-art performance without requiring additional data, auxiliary pre-trained networks, or extensive hyperparameter tuning.
Figure 1: Attention maps of the final visual encoder layer. For the patches shaded in red (denoted with (a) to (d)), the final layer’s attention maps are presented for CLIP [32], SCLIP [38], and our method. We have identified two problematic phenomena in the attention maps of CLIP and SCLIP, stemming from a lack of mechanisms to properly attend to patches’ neighbourhoods. First, as depicted in (a) and (b), attention intensity is sometimes dispersed among distant patches, neglecting the vicinity of a patch. Additionally, adjacent or closely located patches sharing the same real-world category and even similar visual characteristics can have inconsistent attention maps. For instance, while SCLIP generates a quality attention map for patch (d), its attention map for (c) is notably different and fails to focus on the desired object. By explicitly promoting attention to neighbours, our method produces consistent attention maps across adjacent patches.
@@ 2.1 Adapting large-scale vision-language models
The field of deep learning is undergoing a learning paradigm shift with the emergence of vision-language models trained at a large scale. In particular, the Contrastive Language-Image Pre-Training (CLIP) method [32] has shown unprecedented success, mainly due to its remarkable zero-shot and few-shot transfer capabilities on visual recognition tasks, notably in the context of classification. Fueled by its generalization and transfer capabilities, a realm of approaches have arisen to either improve its zero-shot performance by modifying its training on image-text pairs [47, 39, 14], or efficiently adapt to novel tasks with only a few labeled images [50, 15, 48, 36, 18, 34]. Nonetheless, the pre-training is performed at the image level, and therefore, typically, only the class token is trained to capture the global information, which limits the applicability of these approaches for dense prediction tasks.
@@ 2.2 Open-vocabulary semantic segmentation
The outstanding transferability of CLIP in classification has surged to a rapidly growing literature in open-vocabulary semantic segmentation (OVSS) [2, 16, 22, 44, 45, 46, 52, 33, 43, 26, 45, 27], which attempts to segment novel concepts in a given image without explicit supervision for them. These methods can be categorized into three main groups, depending on the level of auxiliary data required for adaptation: fully-supervised, weakly-supervised, and training-free. Fully-supervised OVSS approaches [2, 16, 22, 44, 45, 46, 31] adapt the pre-trained CLIP to semantic segmentation by exploiting a large labeled set containing pixel-wise annotations from a set of classes. The main idea of these approaches is that, by leveraging a large pixel-labeled dataset of an arbitrary limited set of categories, the model can learn to perform segmentation, while maintaining the remarkable transfer capability of CLIP. Nevertheless, the adaptation datasets employed in this family of approaches typically present a high class overlap with the testing images. Instead of requiring pixel-wise masks, weakly-supervised OVSS resorts to additional datasets with image-level tags, typically in the form of image-text pairs [8, 6, 49, 40, 42, 24, 7]. A common strategy is to use these large image-text pair sets as supervision during adaptation, where the categories present in an image are included in the text. The alignment between the visual and textual information is typically performed via a contrastive loss [42], similar to the pre-training of CLIP, which can be further enhanced by integrating additional learning strategies, such as online pixel clustering [24] or multiple contrastive losses between the features from original images and those restored from corrupted versions [6], for example. Nevertheless, a substantially large dataset is still required for adaptation, and classes appearing in each image need to be known in advance, presenting impractical considerations for real-world problems.
@@
Given the points raised above, we focus in this work on training-free OVSS where, ideally, access to additional data for adaptation is not allowed. Nevertheless, as stated earlier, many of the methods rather rely on leveraging auxiliary models pre-trained on large-scale datasets (e.g., ViT [33, 41, 27] or stable diffusion [3, 19, 11]), or containing multiple components whose hyperparameters must be tuned [3, 27]. A different line of work attempts to improve the potential of CLIP in extracting dense visual features by modifying ViT’s self-attention [52, 4, 21]. This can be done, for instance, by directly using the value vectors [52], or by introducing additional pathways computing self-self attentions in parallel to the encoder blocks to aggregate output of several layers using the residual connections [21, 4]. A main difference with this line of work, however, is that we introduce explicit local spatial consistency in the self-attention, a concept that has been disregarded by the existing training-free OVSS literature.
@@
The concurrent approach, SCLIP [38], proposes a self-attention mechanism encouraging each token to attend to itself and the positions sharing similar information. In particular, the authors state that an attention map with maximum values on the diagonal leads to proper localization. Nevertheless, since query and key vectors are not normalized in SCLIP, there is no guarantee that the maximum values of the attention map fall on the diagonal. Hence, if an outlier patch has a key vector with a magnitude much larger than the others, patches attend to the outlier more than themselves. Furthermore, even though in a well-localized model each patch should attend to itself with a high intensity, localization is not limited to self, and the vicinity of each patch should be taken into account as well. This is particularly important in segmentation, where adjacent patches often represent the same class. As apparent in our qualitative analysis (Fig. 1), despite SCLIP improving the spatial localization of attention maps over CLIP, it often generates contrasting attention maps for adjacent patches containing the same real-world object. This demonstrates that in this method, spatial localization in the neighbourhood of a patch does not typically go far beyond the patch itself. Moreover, although for a given patch, SCLIP attends to the patches that share similar query or key vectors, neighbouring patches do not necessarily do. We argue that only attending to patches that share similar vectors (and also self) is suboptimal, which is demonstrated empirically in Fig. 1. Our proposed method can handle these limitations by adding an explicit mechanism that enforces patches to attend to their neighbours, imposing the local spatial consistency required for semantic segmentation.
@@ 3 Preliminaries
In this work, we address the task of open-vocabulary semantic segmentation in a training-free scenario, i.e., with no supervision and without fine-tuning the parameters. Hence, for a given image $\mathbf{X}_{i}\in{\mathbb{R}}^{\Omega_{i}}$, where $\Omega_{i}$ denotes its spatial domain $H\times W\times C$, and a set of arbitrary concepts described by natural language $\mathbf{t}_{j}\in\mathcal{T}$, our goal is to provide a segmentation mask of concepts present in the image. In particular, an image $\mathbf{X}_{i}$ is fed into CLIP’s ViT, which produces a representation of each patch. In parallel, $\mathbf{t}_{j}\;\forall j$ are fed into CLIP’s text encoder and the resulting representations are compared to each patch’s representation using cosine similarity, forming a probability distribution which is converted to a segmentation mask using an $\arg\max$ operation.
@@ 3.1 Background on CLIP
CLIP [32] employs a joint training approach, aligning visual and text modalities into the same feature space. Since most of CLIP-based OVSS methods utilize the vision transformer [12] version of CLIP, we focus on this architecture in what follows. In this model, the visual encoder comprises $L$ sequential blocks, each processing $1+hw$ tokens: the first one representing the [CLS] token, which captures global information, and the subsequent tokens each representing a patch. Thus, given an input image, it is initially partitioned into $hw$ non-overlapping patches ($h=H/P$ and $w=W/P$) each in ${\mathbb{R}}^{P\times P\times C}$, where $(P,P)$ denotes the resolution of each patch. Then, a linear transformation projects the sequence of $1+hw$ tokens from $C$ channels to a $D$-dimensional space, and positional embeddings are added to create the input for the first encoder block. The components of this model are described below. For the sake of simplicity of our method’s formulation in Sec. 4.1, we avoid flattening the 2D grid of patches and do not consider the [CLS] token in the formulations below.
Encoder block.
Each encoder block $l$ receives a sequence of tokens, $\mathbf{Z}^{(l-1)}$, from the preceding block and performs the following operations:
@@
In these equations, $\operatorname{LN}$, $\operatorname{SA}$ and $\operatorname{MLP}$ denote layer normalization, the self-attention module, and the feed-forward neural network, respectively. Please note that skip connections refer to the addition of $\mathbf{Z}^{(l-1)}$ and $\mathbf{Z}^{\prime}$ in Eqs. 2 and 4. Furthermore, the combination of Eqs. 1 and 2 is commonly referred to as the self-attention block, whereas the combination of Eqs. 3 and 4 is called the feed-forward block.
Self-attention module.
Within the self-attention module, the sequence of token representations undergoes a linear transformation $\mathbf{W}^{qkv}$, yielding three sequences of $d$-dimensional vectors: query ($\mathbf{q}$), key ($\mathbf{k}$), and value ($\mathbf{v}$). Subsequently, a similarity measure is calculated between all the tokens. More specifically, for a given patch at position $(i,j)$ the dot product of $\mathbf{q}_{ij}$ and $\mathbf{k}_{mn}$ is computed for all $m\in\{1,2,\dots,h\}$ and $n\in\{1,2,\dots,w\}$. This measure is then scaled by $1/\sqrt{d}$ and passed through a $\mathrm{softmax}$ operation to calculate a weighted sum of $\mathbf{v}$ for the patch. Finally, the output is obtained via a projection using a linear transformation $\mathbf{W}^{o}$. Thus, formally, operation $\operatorname{SA}$ for the patch at position $(i,j)$ can be described as:
@@
Hereafter, we will refer to $\mathrm{softmax}\left(\text{sim}_{ij}\right)$ as the attention map of point $(i,j)$, and $\text{sim}_{ij}$ as its corresponding logits. Note that while in practice a multi-head version of the self-attention is used, for simplicity, we have presented equations considering a single head.